# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ Appendix B Proof of Theorems
$\|\sqrt{\mathbf{v}_{t}}-\sqrt{\mathbf{v}_{t+1}}\|_{\infty}\leq\sqrt{2(1-\beta_{2,t})}.$ (B.1)
The proof is in Appendix C.3.
Lemma B.2.
In Algorithm 1. Under Assumption 4.1 and 4.2, if $1\geq\beta_{1,t+1}\geq 0$, $\forall t$, we have
$\mathbb{E}\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}$ $\leq\beta_{1,t+1}^{2}\mathbb{E}\|\nabla f(\bm{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+2\beta_{1,t+1}^{2}L^{2}\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+2(1-\beta_{1,t+1})^{2}\sigma^{2}$
$\qquad-M_{t+1}$
@@
$\bm{\Delta}_{t}:=\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1}),\quad M_{t+1}:=\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\Big{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Big{)}^{2}.$ (B.2)
$A_{t+1}:=\frac{G_{t+1}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}}{\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}.$ (B.3)
@@
$G_{t+1}$ $:=(1-\beta_{1,t+1})\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\bigg{\rangle}+\beta_{1,t+1}\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},\bm{\varepsilon}_{t}\bigg{\rangle},$ (B.4)
The proof is in Appendix C.4.
Lemma B.3.
In Algorithm 1. With Assuptions 4.2 and 4.3 and $\eta_{t}\leq\rho\cdot(2L)^{-1},\forall t\geq 1$, it holds that
@@
The proof of Lemma B.3 is in Appendix C.5.
Lemma B.4.
In Algorithm 2. With Assuptions 4.2 and 4.3 and $\eta_{t}\leq\min\{(4\lambda)^{-1},\rho\cdot(2L)^{-1}\},\forall t\geq 1$, it holds that
@@
The proof of Lemma B.4 is in Appendix C.6.
Lemma B.5.
@@
B.1 Proof of Theorem 4.4
Proof of Theorem 4.4.
@@
$\Phi_{t+1}-\Phi_{t}$ $=\underbrace{\mathbb{E}[F(\mathbf{x}_{t+1})-F(\mathbf{x}_{t})]}_{\mbox{$I_{1}$}}+\underbrace{\mathbb{E}\Bigg{[}\frac{\rho}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}-\frac{\rho}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Bigg{]}}_{\mbox{$I_{2}$}}.$ (B.5)
For $I_{1}$, we use Lemma B.3 to obtain
$I_{1}\leq\mathbb{E}\Big{[}-\frac{\rho}{2\eta_{t}}\cdot\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\eta_{t}}{\rho}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Big{]}.$ (B.6)
For $I_{2}$, we use Lemma B.2 to obtain
@@
$\leq\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1},$ (B.7)
@@
From Lemma B.5, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}<\eta_{t}$. Choosing $c$ such that $c\geq 32L^{2}\rho^{-2}+1$, we obtain
$\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\leq\frac{\rho}{16L^{2}}\cdot(\eta_{t}-c\eta_{t})\leq-2\eta_{t}\rho^{-1}.$ (B.8)
Bringing (B.8) into (B.7), we arrive at the upper bound for $I_{2}$:
$I_{2}\leq-\frac{2\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (B.9)
Now combining (B.5), (B.6) and (B.9), we derive
@@
$\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{3\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}$
$\leq\Phi_{1}-\Phi_{T+1}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}}\sum_{t=1}^{T}\frac{1}{s+t}-\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}$
@@
$\Phi_{1}$ $=\mathbb{E}\Big{[}F(\mathbf{x}_{1})+\frac{\rho s^{1/3}}{16L^{2}}\cdot\|\nabla F(\mathbf{x}_{1})-\mathbf{m}_{1}\|_{2}^{2}\Big{]}$
$=F(\mathbf{x}_{1})+\frac{\rho s^{1/3}}{16L^{2}}\cdot\mathbb{E}[\|\nabla F(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{1},\bm{\xi}_{1})\|_{2}^{2}]$
$\leq F(\mathbf{x}_{1})+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}.$
@@
$\frac{1}{T}\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{3\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}$ (B.10)
$\leq\frac{G}{T}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}T}\cdot\log(s+T)-\frac{1}{T}\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (B.11)
@@
$\leq\Big{(}2\rho G+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{16L^{2}T^{1/3}},$
where the last inequality holds when $T\geq s$. Similarly, for the second term in (B.11), we also have
@@
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\frac{8G}{3\rho T\eta_{T}}+\frac{c^{2}\sigma^{2}}{3\rho L^{2}T\eta_{T}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{2/3}}$
$\leq\frac{16G}{3\rho T^{2/3}}+\frac{2c^{2}\sigma^{2}}{3\rho L^{2}T^{2/3}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{1/3}}.$
@@
B.2 Proof of Theorem 4.6
Proof of Theorem 4.6.
@@
$\qquad+\underbrace{\mathbb{E}\Bigg{[}\frac{\rho}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}-\frac{\rho}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Bigg{]}}_{\mbox{$I_{2}$}}.$ (B.12)
For $I_{1}$, we use Lemma B.4 to obtain
$I_{1}\leq\mathbb{E}\Bigg{[}-\frac{\rho}{4\eta_{t}}\cdot\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\eta_{t}}{\rho}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\lambda D^{2}\sqrt{2(1-\beta_{2,t})}}{2}\Bigg{]}.$ (B.13)
For $I_{2}$, we use Lemma B.2 to obtain
@@
$\leq\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1},$ (B.14)
@@
From Lemma B.5, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}<\eta_{t}$. Choosing $c$ such that $c\geq 32L^{2}\rho^{-2}+1$, we obtain
$\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\leq\frac{\rho}{16L^{2}}\cdot(\eta_{t}-c\eta_{t})\leq-2\eta_{t}\rho^{-1}.$ (B.15)
Bringing (B.15) into (B.14), we arrive at the upper bound for $I_{2}$:
$I_{2}\leq-\frac{2\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (B.16)
Now combining Formulas (B.12), (B.13) and (B.16), we obtain
@@
where the last inequality follows by taking $\beta_{2,t}=1-\eta_{t}^{6}$. By the definition of $\Phi_{t}$, we have $\Phi_{T+1}\geq F(\mathbf{x}_{t+1})\geq\min_{\mathbf{x}}F(\mathbf{x})$. And for $\Phi_{1}$, according to the fact following (C.7) that $\|\mathbf{H}_{t+1}\|_{2}=\Big{\|}\text{diag}(\sqrt{\mathbf{v}_{t+1}}+\epsilon)\Big{\|}_{2}\leq 1+\epsilon,$ we obtain
@@
$\frac{1}{T}\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}$ (B.17)
$\leq\frac{G}{T}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}T}\cdot\log(s+T)+\frac{\lambda D^{2}\log(s+T)}{T}-\frac{1}{T}\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (B.18)
@@
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\frac{8(G+\lambda D^{2}\log(s+T))}{\rho T\eta_{T}}+\frac{c^{2}\sigma^{2}}{\rho L^{2}T\eta_{T}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{2L^{2}T\eta_{t}^{2}}$
$\qquad\leq\frac{16(G+\lambda D^{2}\log(s+T))}{\rho T^{2/3}}+\frac{2c^{2}\sigma^{2}}{\rho L^{2}T^{2/3}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{2L^{2}T^{1/3}}.$
@@
Appendix C Proof of Auxiliary Lemmas
C.1 Lemma C.1 and Proof
Lemma C.1.

### v3

@@ Appendix C Proof of Theorems
$\|\sqrt{\mathbf{v}_{t}}-\sqrt{\mathbf{v}_{t+1}}\|_{\infty}\leq\sqrt{2(1-\beta_{2,t})}.$ (C.1)
The proof is in Section D.3.
Lemma C.2.
In Algorithm 1. Under Assumption B.1 and B.2, if $1\geq\beta_{1,t+1}\geq 0$, $\forall t$, under approximate choice of $\gamma_{t+1}$ as in (D.12), we have
$\mathbb{E}\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}$ $\leq\beta_{1,t+1}^{2}\mathbb{E}\|\nabla F(\bm{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+2\beta_{1,t+1}^{2}L^{2}\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+2(1-\beta_{1,t+1})^{2}\sigma^{2}-M_{t+1}$
@@
$M_{t+1}:=\mathbb{E}\|\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\|_{2}^{2}\Bigg{(}A_{t+1}^{2}-\Big{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Big{)}^{2}\Bigg{)},$ (C.2)
$A_{t+1}:=\frac{G_{t+1}+\beta_{1,t+1}\text{tr}\big{(}\text{Var}\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\big{)}\big{)}}{\mathbb{E}\|\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\|_{2}^{2}}$
@@
$G_{t+1}$ $:=(1-\beta_{1,t+1})\mathbb{E}\bigg{\langle}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1}),\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\bigg{\rangle}$
$+\beta_{1,t+1}\mathbb{E}\bigg{\langle}\nabla F(\mathbf{x}_{t+})-\nabla F(\mathbf{x}_{t}),F(\mathbf{x}_{t})-\mathbf{m}_{t}\bigg{\rangle}.$
The proof is in Section D.4.
Lemma C.3.
In Algorithm 1. With Assuptions B.2 and B.3 and $\eta_{t}\leq\rho\cdot(2L)^{-1},\forall t\geq 1$, it holds that
@@
The proof of Lemma C.3 is in Section D.5.
Lemma C.4.
In Algorithm 2. With Assuptions B.2 and B.3 and $\eta_{t}\leq\min\{(4\lambda)^{-1},\rho\cdot(2L)^{-1}\},\forall t\geq 1$, it holds that
@@
The proof of Lemma C.4 is in Section D.5.
Lemma C.5.
@@
C.1 Proof of Theorem B.5
Proof of Theorem B.5.
@@
$\Phi_{t+1}-\Phi_{t}$ $=\underbrace{\mathbb{E}[F(\mathbf{x}_{t+1})-F(\mathbf{x}_{t})]}_{\mbox{$I_{1}$}}+\underbrace{\mathbb{E}\Bigg{[}\frac{\rho}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}-\frac{\rho}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Bigg{]}}_{\mbox{$I_{2}$}}.$ (C.3)
For $I_{1}$, we use Lemma C.3 to obtain
$I_{1}\leq\mathbb{E}\Big{[}-\frac{\rho}{2\eta_{t}}\cdot\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\eta_{t}}{\rho}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Big{]}.$ (C.4)
For $I_{2}$, we use Lemma C.2 to obtain
@@
$\leq\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1},$ (C.5)
@@
From Lemma C.5, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}<\eta_{t}$. Choosing $c$ such that $c\geq 32L^{2}\rho^{-2}+1$, we obtain
$\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\leq\frac{\rho}{16L^{2}}\cdot(\eta_{t}-c\eta_{t})\leq-2\eta_{t}\rho^{-1}.$ (C.6)
Bringing (C.6) into (C.5), we arrive at the upper bound for $I_{2}$:
$I_{2}\leq-\frac{2\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (C.7)
Now combining (C.3), (C.4) and (C.7), we derive
@@
$\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{3\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}$ $\leq\Phi_{1}-\Phi_{T+1}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}}\sum_{t=1}^{T}\frac{1}{s+t}-\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}$
@@
$\Phi_{1}$ $=\mathbb{E}\Big{[}F(\mathbf{x}_{1})+\frac{\rho s^{1/3}}{16L^{2}}\cdot\|\nabla F(\mathbf{x}_{1})-\mathbf{m}_{1}\|_{2}^{2}\Big{]}=F(\mathbf{x}_{1})+\frac{\rho s^{1/3}}{16L^{2}}\cdot\mathbb{E}[\|\nabla F(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{1},\bm{\xi}_{1})\|_{2}^{2}]\leq F(\mathbf{x}_{1})+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}.$
@@
$\frac{1}{T}\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{3\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}\leq\frac{G}{T}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}T}\cdot\log(s+T)-\frac{1}{T}\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (C.8)
@@
$\leq\Big{(}2\rho G+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{8L^{2}T^{1/3}},$
where the last inequality holds when $T\geq s$. Similarly, for the second term in (C.8), we also have
@@
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\frac{8G}{3\rho T\eta_{T}}+\frac{c^{2}\sigma^{2}}{3L^{2}T\eta_{T}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{2/3}}$
$\leq\frac{16G}{3\rho T^{2/3}}+\frac{2c^{2}\sigma^{2}}{3L^{2}T^{2/3}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{1/3}}.$
@@
C.2 Proof of Theorem B.6
Proof of Theorem B.6.
@@
$+\underbrace{\mathbb{E}\Bigg{[}\frac{\rho}{16L^{2}\eta_{t}}\cdot\|\nabla F(\mathbf{x}_{t+1})-\mathbf{m}_{t+1}\|_{2}^{2}-\frac{\rho}{16L^{2}\eta_{t-1}}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}\Bigg{]}}_{\mbox{$I_{2}$}}.$ (C.9)
For $I_{1}$, we use Lemma C.4 to obtain
$I_{1}\leq\mathbb{E}\Bigg{[}-\frac{\rho}{4\eta_{t}}\cdot\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\eta_{t}}{\rho}\cdot\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\lambda D^{2}\sqrt{2(1-\beta_{2,t})}}{2}\Bigg{]}.$ (C.10)
For $I_{2}$, we use Lemma C.2 to obtain
@@
$\leq\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1},$ (C.11)
@@
From Lemma C.5, we know that $\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}<\eta_{t}$. Choosing $c$ such that $c\geq 32L^{2}\rho^{-2}+1$, we obtain
$\frac{\rho}{16L^{2}}\cdot\Big{(}\frac{\beta_{1,t+1}^{2}}{\eta_{t}}-\frac{1}{\eta_{t-1}}\Big{)}\leq\frac{\rho}{16L^{2}}\cdot(\eta_{t}-c\eta_{t})\leq-2\eta_{t}\rho^{-1}.$ (C.12)
Bringing (C.12) into (C.11), we arrive at the upper bound for $I_{2}$:
$I_{2}\leq-\frac{2\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\frac{\rho c^{2}\eta_{t}^{3}\sigma^{2}}{8L^{2}}-\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (C.13)
Now combining Formulas (C.9), (C.10) and (C.13), we obtain
@@
where the last inequality follows by taking $\beta_{2,t}=1-\eta_{t}^{6}$. By the definition of $\Phi_{t}$, we have $\Phi_{T+1}\geq F(\mathbf{x}_{t+1})\geq\min_{\mathbf{x}}F(\mathbf{x})$. And for $\Phi_{1}$, according to the fact following (D.7) that $\|\mathbf{H}_{t+1}\|_{2}=\Big{\|}\text{diag}(\sqrt{\mathbf{v}_{t+1}}+\epsilon)\Big{\|}_{2}\leq 1+\epsilon,$ we obtain
@@
$\frac{1}{T}\sum_{t=1}^{T}\Big{(}\frac{\eta_{t}}{\rho}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}+\frac{\rho}{8\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}\Big{)}$ (C.14)
$\leq\frac{G}{T}+\frac{\rho c^{2}\sigma^{2}}{8L^{2}T}\cdot\log(s+T)+\frac{\lambda D^{2}\log(s+T)}{T}-\frac{1}{T}\sum_{t=1}^{T}\frac{\rho}{16L^{2}\eta_{t}}M_{t+1}.$ (C.15)
@@
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\frac{8(G+\lambda D^{2}\log(s+T))}{\rho T\eta_{T}}+\frac{c^{2}\sigma^{2}}{L^{2}T\eta_{T}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{2L^{2}T\eta_{t}^{2}}$
$\qquad\leq\frac{16(G+\lambda D^{2}\log(s+T))}{\rho T^{2/3}}+\frac{2c^{2}\sigma^{2}}{L^{2}T^{2/3}}\cdot\log(s+T)-\frac{\sum_{t=1}^{T}M_{t+1}}{L^{2}T^{1/3}}.$
@@
Appendix D Proof of Auxiliary Lemmas
D.1 Lemma D.1 and Proof
Lemma D.1.